OpenStudy (anonymous):

given sinx=-4/5 and x is in quadrant 3, what is the value of tan(x/2)

OpenStudy (mertsj):



OpenStudy (mertsj):

\[\tan\frac{x}{2}=\frac{\sin x}{1+\cos x}=\frac{\frac{-4}{5}}{1+\frac{-3}{5}}=-2\]